OpenStudy (katherinesmith):

what is the approximate value of the function at x = -4? there is a graph.

OpenStudy (katherinesmith):

this is the graph

zepdrix (zepdrix):

So when we look over at x=-4, we try to figure out at what height the red line is located. Looks like it's between 1 and 2 yes? Maybe 1.75ish?

OpenStudy (katherinesmith):

is it closer to 1.5 or 1.8? i'm assuming 1.8

zepdrix (zepdrix):

Yah 1.8 sounds a little bit more accurate :)

zepdrix (zepdrix):

It's probably 1.5 at around x=-6

OpenStudy (katherinesmith):

i have another one so stay here!

OpenStudy (katherinesmith):

what is the approximate value of the function at x = 2?

zepdrix (zepdrix):

Hmm, so our red line is below the x-axis at x=2, right? Take a guess! :)

OpenStudy (katherinesmith):

either 0.25 or 0.50... positive or negative

zepdrix (zepdrix):

Since it's below the x-axis, it will definitely be negative. It's between -1 and 0. Yah -0.25 seems like a good guess!

OpenStudy (katherinesmith):

do you know how to do ones like : solve x/(x-2) < 2

zepdrix (zepdrix):

\[\Large \frac{x}{x-2}\lt2\]Let's ummmm, multiply both sides by x-2,\[\Large x \lt 2(x-2)\]Understand that step? :o The (x-2)'s cancel on the left side.

OpenStudy (katherinesmith):

yes.

OpenStudy (katherinesmith):

so that leaves you with x < 2x - 4

OpenStudy (katherinesmith):

then you put x on one side and get -x < - 4, then it becomes x > 4. correct answer?

zepdrix (zepdrix):

Hmm one sec, yes that's part of our inequality. I think we should be getting another part also though. Lemme see what I messed up.

OpenStudy (katherinesmith):

x > 4 could be a correct answer... i don't know why it wouldn't be. but let me know what you get.

zepdrix (zepdrix):

Oh you have some multiple choice answers or no? :

OpenStudy (katherinesmith):

no i don't.

zepdrix (zepdrix):

oh :)

zepdrix (zepdrix):

The problem is this... Remember what happens when we multiply/divide across an inequality? The inequality sign `flips` directions. Example:\[\Large -3x \lt 6\]Dividing both sides by -3 gives us,\[\Large x>-2\]See how the sign changed from `less than` to `greater than` ?

OpenStudy (katherinesmith):

yes, which i did that.

zepdrix (zepdrix):

So in our problem, I think it's a bad idea to multiply both sides by (x-2) as I did. We don't know whether or not that quantity is negative. So let's try something else :O

OpenStudy (katherinesmith):

okay

zepdrix (zepdrix):

\[\Large \frac{x}{x-2}\lt2\]Subtracting 2 from each side,\[\Large \frac{x}{x-2}-2<0\]Let's get a common denominator:\[\Large \frac{x}{x-2}-\frac{2(x-2)}{x-2}\lt0\]

zepdrix (zepdrix):

So what does that simplify down to? :)\[\Large \frac{x-2(x-2)}{x-2}<0\]

zepdrix (zepdrix):

Hmm I hope I'm doing this right, man I hate inequalities lol. Lemme check one of the smarty pants guys :P @SithsAndGiggles @satellite73

OpenStudy (katherinesmith):

let me know. im still confused

OpenStudy (anonymous):

I don't think putting the 2 on the other side is really necessary. I'd go with \[x<2(x-2)\] which gives \[x<2x-4\\ -x<-4\] Either way works, but I think this is simpler. ;)

OpenStudy (katherinesmith):

so that leaves you with x > 4. because you divide by -x and flip the inequality.

zepdrix (zepdrix):

Well if you wolfram this, it gives 2 intervals, x<2, x>4. So I was just getting a little confused :d

OpenStudy (katherinesmith):

how did you get the two intervals? because thats' the answer i need

zepdrix (zepdrix):

Because from this point, where I combined the fractions:\[\Large \frac{x-2(x-2)}{x-2}<0\]After we simplify, we get,\[\Large \frac{-x+4}{x-2}\lt0\]

zepdrix (zepdrix):

If we look at the top and bottom separately (which I guess is what we're supposed to do),\[\Large -x+4\lt0 \qquad\qquad\qquad x-2\lt0\]

zepdrix (zepdrix):

\[\Large -x+4\lt0 \qquad\to\qquad x>4\]\[\Large x-2\lt0 \quad\qquad\to\qquad x<2\]

zepdrix (zepdrix):

That too confusing? :d Or was it back when we combined the fractions the part you got stuck on?

OpenStudy (katherinesmith):

this all just hurts my head. lord jesus help me

zepdrix (zepdrix):

XD

OpenStudy (katherinesmith):

alright one more question

zepdrix (zepdrix):

k c:

OpenStudy (katherinesmith):



OpenStudy (katherinesmith):

that's an = sign and then a - sign

zepdrix (zepdrix):

Solve for x?

OpenStudy (katherinesmith):

yes

zepdrix (zepdrix):

Note:\[\Large x^2+x \qquad=\qquad x(x+1)\]

zepdrix (zepdrix):

\[\Large \frac{x+5}{x(x+1)}=\frac{1}{x(x+1)}-\frac{x-6}{x+1}\]

OpenStudy (katherinesmith):

then what the heck do you do

zepdrix (zepdrix):

If we multiply both sides by (x+1), can you see what will happen? :o

OpenStudy (katherinesmith):

please write it out for me

zepdrix (zepdrix):

Lemme give you a simpler example really quick :) Maybe it will help it to click.\[\Large \frac{x}{3}=\frac{7}{3}\] If we multiply both sides by 3, what does it give us?

OpenStudy (katherinesmith):

x = 7

zepdrix (zepdrix):

good, all of the denominator 3's cancel with the 3's that we multiplied through by, right?

OpenStudy (katherinesmith):

yes

OpenStudy (katherinesmith):

\[\frac{ x + 5 }{ x } = \frac{ 1 }{ x } - x - 6\]

OpenStudy (katherinesmith):

then what

zepdrix (zepdrix):

One small thing to be careful of. We don't usually write it, but there is a pair of parenthesis implied when we have a negative in front of a fraction.\[\Large \frac{x+5}{x(x+1)}=\frac{1}{x(x+1)}-\frac{(x-6)}{x+1}\]

zepdrix (zepdrix):

\[\Large \frac{ x + 5 }{ x } = \frac{ 1 }{ x } -(x - 6)\]

OpenStudy (katherinesmith):

I got that. what next.

zepdrix (zepdrix):

To finish getting rid of the fractions, let's multiply both sides by x.

zepdrix (zepdrix):

Gives us something like this, yes? :)\[\Large (x+5)=1-x(x-6)\]

OpenStudy (katherinesmith):

yes

zepdrix (zepdrix):

From here, distribute your -x to each term in the brackets. And then try to get everything to one side, so the other side =0

zepdrix (zepdrix):

crap i gotta go :U sorryyyy :C gotta go take care of something